# Motion-Zero: Zero-Shot Moving Object Control Framework for Diffusion-Based Video Generation

### v3

@@ Abstract
Recent large-scale pre-trained diffusion models have demonstrated a powerful generative ability to produce high-quality videos from detailed text descriptions. However, exerting control over the motion of objects in videos generated by any video diffusion model is a challenging problem. In this paper, we propose a novel zero-shot moving object trajectory control framework, Motion-Zero, to enable a bounding-box-trajectories-controlled text-to-video diffusion model. To this end, an initial noise prior module is designed to provide a position-based prior to improve the stability of the appearance of the moving object and the accuracy of position. In addition, based on the attention map of the U-net, spatial constraints are directly applied to the denoising process of diffusion models, which further ensures the positional and spatial consistency of moving objects during the inference. Furthermore, temporal consistency is guaranteed with a proposed shift temporal attention mechanism. Our method can be flexibly applied to various state-of-the-art video diffusion models without any training process. Extensive experiments demonstrate our proposed method can control the motion trajectories of objects and generate high-quality videos.
Figure 1: Our Motion-Zero framework endows any pre-trained video diffusion model with the capability to manipulate object trajectories directly, circumventing the need for supplementary training. By designating the target entity in the input prompts and a sequence of bounding boxes, users can intuitively direct the motion path of the object within the generated video sequence.
1 Introduction
@@
In recent years, the generative capabilities of diffusion models have been widely recognized in both text-to-image Dhariwal and Nichol (2021); Ho et al. (2020); Rombach et al. (2022); Song et al. (2020) and text-to-video domains Guo et al. (2023); Ho et al. (2022). Although these video models are capable of producing high-definition, high-resolution, and fluid video animations, the dynamic motion trajectories of the generated objects are relatively random in existing text-to-video models Blattmann et al. (2023); Zhang et al. (2023a). Thus, accurate control of an object’s motion trajectories in a generated video remains rudimentary.
@@
To address this problem, various efforts have been proposed in different strategies. Some methods Zhang et al. (2023b); Wu et al. (2023a) try to control the trajectory of the moving object by providing a detailed conditional motion sequence, such as a dancing skeleton. However, the cost of acquiring conditional control sequences is non-negligible and inhibits the diverse generation of video outputs by the user. To obtain variety over the generated video with more control capability, i.e. trajectories of moving objects, several methods Wang et al. (2023c, b); Yin et al. (2023) harness a substantial dataset of motion and trajectory pairs to train the baseline models. Although the results are very impressive and promising, these methods are solely applicable to the base models on which they were trained and cannot be applied to other models directly. Furthermore, all of these methods require extensive training and significant computational resources, preventing ordinary users from using them.
@@
Nevertheless, the fundamental text-to-video diffusion models Blattmann et al. (2023); Zhang et al. (2023a) have undergone an extremely large scale training dataset Bain et al. (2021); Xue et al. (2022). Theoretically, the pre-trained model should inherently have learned extensive knowledge about the dynamics of a variety of object movements. However, the semantics of the latent space in different modules of a video diffusion model are not explicitly defined, which makes it difficult to manually control generated video in desired motion dynamics. To some extent, methods mentioned above try to align and define those latent spaces in the baseline model by training with additional labeled datasets, and successfully harness the intrinsic knowledge of pre-trained models to control the motion. In addition, we observe that the initial noise plays a significant inspirational role in the generation of the videos. The same initial noise tends to produce videos with similar content. With these two characteristics, we can manipulate the trajectory of moving objects.
@@
In this paper, we propose a zero-shot framework, Motion-Zero, that can be easily applied to any pre-trained video diffusion model and achieve universality and plug-and-play capabilities of controlling. To this end, we designed an initial noise prior module to provide noises based on the motion trajectories given by the user. Inspired by the fact that the values of the cross-attention map within the U-Net largely determine the generation location of the subject mentioned in the prompt Xie et al. (2023); Epstein et al. (2023), spatial constraints with an attention similarity loss over these attention maps are proposed to achieve precise manipulation of object positioning within individual frames. Furthermore, based on our observations, we find that merely imposing spatial constraints on the positions of objects can negatively impact the overall coherence and video quality. To preserve the continuity throughout the video sequence, a novel shift temporal attention mechanism is proposed to further maintain the baseline model focus on the same objects across the time axis. Through the implementation of these modules, our framework can effectively generate high-quality video controlled by bounding-box trajectories without any training process. Samples generated by Motion-Zero can be viewed in Fig.1. The main innovations of our work can be summarized as follows,
@@
  • We propose a zero-shot framework Motion-Zero which is capable of controlling the motion trajectories of arbitrary objects within a pre-trained video generation diffusion model. Our Motion-Zero is plug-and-play and without any additional training.
  • By recognizing the influence of initial noise on video generation, we implemented an initial noise conditioning module that establishes advantageous starting conditions for controllable video diffusion.
  • To further accurate control over the generative video, we propose spatial constraints and a novel shift temporal attention mechanism aimed at ensuring the positional accuracy and temporal continuity and consistency of the target, respectively.
@@
  • Extensive experiments demonstrate that our proposed work can be applied to any pre-trained video diffusion models. Our results can outperform those from extensively pre-trained trajectory control models.
2 Related Works
2.1 Text-To-Video Diffusion Models
As text-to-image diffusion models have flourished and demonstrated impressive results, the body of work related to text-to-video diffusion has gradually expanded. Researchers and developers are leveraging the principles of text-to-image models to extend their capabilities into the dynamic and temporally complex domain of video generation. In the realm of text-to-video (T2V) synthesis, the field has witnessed several pioneering approaches. Initial endeavors such as Imagen Video Ho et al. (2022) and Make-A-Video Singer et al. (2022) tackled T2V generation at the pixel level, leading to notable advancements but also encountering constraints in video duration and quality due to computational intensity. To address these limitations, MagicVideo Zhou et al. (2022) emerged, introducing a novel autoencoder trained specifically on video data, which, akin to the impact of Latent Diffusion Models (LDMs) Rombach et al. (2022) in the text-to-image (T2I) sector, significantly improved the computational efficiency of T2V tasks. In the current landscape, there exists an abundance of generalizable open-domain text-to-video generation models capable of producing high-definition videos with dynamic visual fidelity Blattmann et al. (2023); Zhang et al. (2023a); Mullan et al. (2023); Sterling (2023); Chen et al. (2023); Wang et al. (2023a). The models often generate either stochastic motion paths or demonstrate relatively minor variations in dynamics, underscoring the potential benefits of investigating methodologies for manually stipulating more deliberate and varied motion trajectories.
2.2 Controllable Video Generalization
@@
There are now pre-trained methodologies Wang et al. (2023c, b); Yin et al. (2023) in place that have the capability to control the motion trajectories of objects within generated videos. VideoComposer Wang et al. (2023b) employs a two-stage training strategy to incrementally incorporate temporal information and control signals. A motion condition encoder is proposed for training, designed to enhance the model’s ability to understand and integrate motion-specific information. DragNUMA Yin et al. (2023) also employs a comparable training strategy and utilizes optical flow as a conditioning mechanism for trajectory modeling. MotionCtrl Wang et al. (2023c) synthesizes the approaches of these two works and introduces a camera control module and a trajectory control module to further refine the management of movement within generated video content. These models typically rely on extensive training on large-scale datasets such as WebVid-10M Bain et al. (2021) and HD-VILA-100M Xue et al. (2022) which lead to expensive training costs. Also, these models are typically constrained to operate on the specific models they were trained on, lacking the flexibility to interchange base video diffusion models. This limitation hinders their adaptability and limits the scope of their application to only those scenarios for which they were explicitly designed. Our model stands out in that it can be applied to any base diffusion model without the necessity for further training. This versatility not only broadens its applicability but also results in substantial savings in training costs and resources, offering a significant advantage over more rigid, model-specific methods. A contemporaneous work Ma et al. (2023) parallels our approach with a similar zero-shot setting. However, they ignore the importance of the inital noise and the control of spatial and temporal consistency.
Figure 2: Overview of our Motion-Zero. The total pipeline is shown on (a). Given the box condition $\mathcal{B}$ and the prompt condition, we generate the prior latents $\mathbf{z}_{T}$ by our Initial Noise Prior Module (INPM) as shown on (b). At timestep $t$, $\mathbf{z}_{t}$ is firstly optimized to $\mathbf{z}_{t}^{\prime}$ by the Spatial Constraints (SC). Subsequently, $\mathbf{z}_{t}^{\prime}$ is passed to the Unet with Shift Temporal Attention Module (STAM) as demonstrated on (c). All the parameters of the video diffusion are frozen. $T_{1}$ represents the number of timesteps during which SC and STAM are applied, and $T_{2}$ denotes the number of timesteps where the original video diffusion process is utilized.
3 Preliminaries
@@
Video Diffusion Model: Video diffusion models are designed to produce high-quality and diverse videos, guided by text prompts. To save the computational costs, Rombach et al. (2022) utilize a U-Net as a denoising model within a latent space, significantly reducing the computational load in terms of both time and space.
@@
Specifically, users can input a text condition $\mathbf{c}$ along with a sequence of rectangular boxes =$\mathcal{B}=\{(x_{1}^{f},y_{1}^{f}),(x_{2}^{f},y_{2}^{f})\}^{N_{f}}$, where $(x_{1}^{f},y_{1}^{f}),(x_{2}^{f},y_{2}^{f})$ are the upper left and lower right points of the box in frame $f$, $N_{f}$ is the total frame number. The boxes $\mathcal{B}$ correspond to an the position of the moving object.
4 Methodology
4.1 Overview
@@
Existing trajectory control methods require large-scale video training data to optimize Eq.3 leading to a high computational cost. Differently, our work falls into the setting of zero-shot motion trajectory control. To this end, our Motion-Zero framework operates entirely during the inference stage, thus eliminating the need for any training and making it compatible with any pre-trained video diffusion model.
In the following section, we provide a detailed presentation of the proposed Motion-Zero framework and its individual components. The pipeline is shown in Fig.2 (a). Sec.4.2 introduces the designed initial noise prior module. Sec.4.3 is followed to describe how we constrain the moving object position and spatial consistency. Finally, Sec.4.4 details our shift temporal attention mechanism which improves the temporal consistency.
4.2 Initial Noise Prior Module
@@
According to the theory of DDIM Inversion Ho et al. (2020), the initial noise has a significant impact on the final generated outcome. The Initial Noise Prior Module (INPM) is designed to leverage this property to provide a strong prior for the position of the moving object. Several steps are involved to integrate a moving object into a sequence of frames with a coherent prior as shown in Fig.2 (b). Firstly, a meta video $V_{meta}$ is sampled using the original Video Diffusion with spatial constraints (introduced in Sec.4.3), $\mathbf{z}^{*}\sim\mathcal{N}(0,\mathbf{I})$ as latent input, conditioned on a given prompt $\mathbf{c}$ and the first frame box $\mathcal{B}^{0}$. Then, a video latent $\mathbf{z}_{meta}$ is generated based on $V_{meta}$ from Encoder $\mathcal{E}$. Once $\mathbf{z}_{meta}$ is prepared, we perform a DDIM Inversion to obtain the corresponding noise latent representation $\mathbf{z}_{I}$. We crop the latent representation within the box $\mathcal{B}^{0}$ for each frame, creating a sequence of latent patches. Subsequently, we use a local mixup operation Zhang et al. (2017) to mix the latent patches and the initial noise $\mathbf{z}^{*}$ in the range of $\mathcal{B}^{f}$ frame by frame. This method allows us to set a coherent prior in the corresponding object’s position in the initial noises. It also ensures that the animated object maintains consistency in appearance and movement across the video frames, without incurring additional computational costs during the generation process. Details are shown in Alg.1.
@@
Function: $\mathbf{Pixel}(\mathbf{a},\mathbf{b})$ means the elements get from tensor $\mathbf{a}$ in the range of box $\mathbf{b}$
Output: Initial noise $\mathbf{z}^{*}$
@@
4.3 Spatial Constraints with Attention Map
The INPM alone is insufficient for precise manipulation of an object’s trajectory. To further improve the capability of control, we introduce Spatial Constraints (SC) deployed at each denoising step $t$ to optimize the intermediate latent representation $\mathbf{z}_{t}$. This optimization is crucial for enhancing the accuracy of the moving object position and preserving spatial consistency. Within the conditional denoising architecture, cross-attention serves as the pivotal bridge that connects the text prompt with the content generated. During the denoising steps, conditioned on the prompt $\mathbf{c}$ and the intermediate features $\mathbf{z}_{t}$, the corresponding cross-attention map can be abtained as $\mathbf{A}$:
@@ 4.3 Spatial Constraints with Attention Map
where $\mathbf{Q},\mathbf{K}$ are the query and key of the $\mathbf{z}_{t}$ and $\mathbf{c}$, respectively. $\mathbf{W}_{Q},\mathbf{W}_{K}$ are two learnable matrices, which are frozen in our settings. Assuming a maximum number $N_{p}$ of prompt tokens $\{\mathbf{p}_{1},...,\mathbf{p}_{N_{p}}\}$ in condition $\mathbf{c}$, at time step $t$, it is feasible to derive $N_{p}$ cross-attention maps $\{\mathbf{A}_{1},...,\mathbf{A}_{N_{p}}\}$.
@@
$[W_{\mathbf{A}_{k}^{f}},H_{\mathbf{A}_{k}^{f}}]=\frac{1}{\sum_{w,h}\mathbf{A}_{k,w,h}^{f}}[\sum_{w,h}w\cdot\mathbf{A}_{k,w,h}^{f},\sum_{w,h}h\cdot\mathbf{A}_{k,w,h}^{f}],$
${\mathcal{L}_{c}}^{f}=||(\frac{x_{1}^{f}+x_{2}^{f}}{2},\frac{y_{1}^{f}+y_{2}^{f}}{2})-(W_{\mathbf{A}_{k}^{f}},H_{\mathbf{A}_{k}^{f}})||_{1},$
where $\mathbf{A}_{k,w,h}^{f}$ is the value of $\mathbf{A}_{k}^{f}$ at the position of $(w,h)$, $(W_{\mathbf{A}_{k}^{f}},H_{\mathbf{A}_{k}^{f})}$ is the centroid position of $\mathbf{A}_{k}^{f}$. After establishing robust control over the position of the object, we recognize unexpected variations in the object’s appearance due to the substantial extent of movement. To ensure appearance consistency across frames, we strive to maintain the uniformity of the $\mathbf{A}_{k}^{f}$ within the box. To this end, we introduce a similarity loss $\mathcal{L}_{s}$ as
${\mathcal{L}_{s}}=1-\frac{1}{N_{f}-1}\sum_{f}^{N_{f}-1}\mathbf{Sim}(\mathbf{Pixel}(\mathbf{A}_{i}^{f},\mathcal{B}^{f}),\mathbf{Pixel}(\mathbf{A}_{i}^{f+1},\mathcal{B}^{f+1})),$
where $\mathbf{Sim}(\cdot,\cdot)$ means the similarity of two elements and $\mathbf{Pixel}()$ is a function getting corresponding elements of the map $\mathbf{A}^{f}$ within a box range $\mathcal{B}^{f}$. Cosine similarity is adopted here. At each timestep, the overall spatial constraints $\mathcal{L}_{sp}$ are formulated as follows:
${\mathcal{L}_{sp}}=\sum_{f}(\lambda_{i}{{\mathcal{L}_{i}}^{f}+\lambda_{o}\mathcal{L}_{o}}^{f}+\lambda_{c}{\mathcal{L}_{c}}^{f})+\lambda_{s}{\mathcal{L}_{s}},$
where $\lambda_{i},\lambda_{o},\lambda_{c},\lambda_{s}$ are hyper-parameters. By minimizing and calculating the gradient of the Eq.11, we can optimize our latent $\mathbf{z}_{t}$ in Eq.2 as follows: S
@@
4.4 Shift Temporal Attention Mechanism
After the spatial constraints on attention maps, pre-trained video diffusion models still encounter difficulties in generating sequences of continuous actions. Within the temporal module of the diffusion process, the latent representation is reshaped into the following configuration (using einops Rogozhnikov (2021) notation):
[…]
To overcome this inconsistency, we propose a Shift Temporal Attention Mechanism (STAM) to improve the dynamics of the moving object in different frames. Specifically, we shift the elements of $\mathbf{z}^{f}$ inside the $\mathcal{B}^{f}$ range with the elements inside the $\mathcal{B}^{0}$ range, as shown in Fig.2 (b). Therefore, the subsequent frames within the box range can be aligned with the box range of the first frame. The steps are shown as follows,
@@ 4.4 Shift Temporal Attention Mechanism
where $\mathbf{Shift}(\cdot,a,b)$ is the shift operation, ${\mathbf{z}^{f}_{w}}$ means the shifted latent ${\mathbf{z}^{f}}$ of frame $f$ and ${\mathbf{z}_{w}}=[{\mathbf{z}^{1}_{w}},..,{\mathbf{z}^{f}_{w}}]$.By applying STAM to a $\mathbf{TemporalAttention}$ layer, we can achieve coherence in the motion of moving objects without additional training, and without incurring extra computational costs during inference.
5 Experiments
5.1 Implementation Details
Figure 3: Quality comparison results on different methods. We take one frame from every three frames. The input prompt of left part: A fish is swimming in the sea. Right part: A bear is walking on the grass. We employed ModelScope (a) and ZeroScope (b) as our baseline models and compared the effect of incorporating additional prompts with the integration of our Motion-Zero. In addition, we conducted a comparative analysis with the pre-trained MotionCtrl model (c). The red dots serve as the control conditions. Zoom in for the best view.

### v4

# Motion-Zero: A Zero-Shot Trajectory Control Framework of Moving Object for Diffusion-Based Video Generation

@@ Abstract
Recent large-scale pre-trained diffusion models have demonstrated a powerful generative ability to produce high-quality videos from detailed text descriptions. However, exerting control over the motion of objects in videos generated by any video diffusion model remains a challenging problem. In this paper, we propose a novel zero-shot moving object trajectory control framework, Motion-Zero, to enable arbitrary single-object-trajectory control for the text-to-video diffusion model. To this end, an initial noise prior module is designed to provide a position-based prior to improve the stability of the appearance of the moving object and the accuracy of position. In addition, based on the attention map of the U-Net, spatial constraints are directly applied to the denoising process of diffusion models, which further ensures the positional consistency of moving objects during the inference. Furthermore, temporal consistency is guaranteed with a proposed shift temporal attention mechanism. Our method can be flexibly applied to various state-of-the-art video diffusion models without any training process. Extensive experiments demonstrate our proposed method can control the motion trajectories of arbitrary objects while preserving the original ability to generate high-quality videos.
Introduction
@@
In recent years, the generative capabilities of diffusion models have been widely recognized in both text-to-image (Dhariwal and Nichol 2021; Ho, Jain, and Abbeel 2020; Rombach et al. 2022; Song, Meng, and Ermon 2020) and text-to-video domains (Guo et al. 2023; Ho et al. 2022). Although these video models are capable of producing high-definition, high-resolution, and fluid video animations, the dynamic motion trajectories of the generated objects are relatively random in existing text-to-video models (Blattmann et al. 2023; Zhang et al. 2023). Thus, accurate control of an object’s motion trajectories in a generated video remains rudimentary.
Figure 1: Our Motion-Zero framework endows different pre-trained video diffusion models with the capability to manipulate object trajectories directly, circumventing the need for supplementary training. By designating the target entity in the input prompts and a sequence of bounding boxes, users can intuitively direct the motion path of the object within the generated video.
@@
Various strategies have been proposed to address this problem. Some methods (Zhang, Rao, and Agrawala 2023; Wu et al. 2023a) try to control the trajectory of the moving object by providing a detailed conditional motion sequence, such as a dancing skeleton. However, the cost of acquiring conditional control sequences is non-negligible and inhibits the diverse generation of video outputs by the user. To obtain variety over the generated video with more control capability, i.e. trajectories of moving objects, several methods (Wang et al. 2023d, c; Yin et al. 2023) harness a substantial dataset of motion and trajectory pairs to train the baseline models. Although the results are very impressive and promising, these methods are solely applicable to the base models on which they were trained and cannot be applied to other models directly. Furthermore, all of these methods require extensive training and significant computational resources, preventing ordinary users from using them.
@@
Nevertheless, the fundamental text-to-video diffusion models (Blattmann et al. 2023; Zhang et al. 2023) have undergone an extremely large scale training dataset (Bain et al. 2021; Xue et al. 2022). Theoretically, the pre-trained model should inherently have learned extensive knowledge about the dynamics of a variety of object movements. However, the semantics of the latent space in different modules of a video diffusion model are not explicitly defined, which makes it difficult to manually control generated video in desired motion dynamics. To some extent, the methods mentioned above try to align and define those latent spaces in the baseline model by training with additional labeled datasets, and successfully harnessing the intrinsic knowledge of pre-trained models to control the motion. In addition, we observe that the initial noise plays a significant inspirational role in the generation of the videos. The same initial noise tends to produce videos with similar content. With these two characteristics, we can manipulate the trajectory of moving objects.
@@
In this paper, we propose a zero-shot trajectory control framework, Motion-Zero, to guide the motion of generated objects in video. Our proposed method can be easily applied to pre-trained video diffusion model while achieving universality and plug-and-play capabilities of controlling. To this end, we first design an Initial Noise Prior Module to provide noises based on the motion trajectories given by the user. Then, inspired by the fact that the values of the cross-attention map within the U-Net largely determine the generation location of the subject mentioned in the prompt (Xie et al. 2023; Epstein et al. 2023), Spatial Constraints with an attention similarity loss over these attention maps are proposed to achieve precise manipulation of object positioning within individual frames. Furthermore, we observe that merely imposing spatial constraints on the positions of objects can negatively impact the overall coherence and video quality. To preserve the continuity throughout the video sequence, a novel Shift Temporal Attention Mechanism is proposed to further maintain the baseline model focus on the same objects across the time axis. Through the implementation of these modules, our proposed framework can effectively generate high-quality video controlled by bounding-box trajectories without any training process. Samples generated by Motion-Zero can be viewed in Fig.1. The main innovations of our work can be summarized as follows,
@@
  • We propose a zero-shot framework Motion-Zero, which is capable of controlling generated object for arbitrary trajectory within a pre-trained video generation diffusion model. Our Motion-Zero is plug-and-play and without any additional training.
  • Our Initial Noise Prior Module is a novel way to acquire semantic initial noise for high-quality generation. Moreover, Spatial Constraints and Shift Temporal Attention Mechanism respectively exploit the cross-attention and temporal-attention dimensions to obtain spatial control and temporal consistency.
@@
  • Extensive experiments demonstrate that our proposed work can be applied to pre-trained video diffusion models. Our method enables the original baseline to generate objects with arbitrary trajectories.
Related Works
@@
With the rapid development of deep learning (Chen et al. 2023b; Zheng et al. 2024b, a; Wang et al. 2023a; Sun et al. 2024b, 2023; Li et al. 2024b, a; Wang et al. 2024c, b; Yang, Li, and Chen 2024; Cai et al. 2021, 2022, 2023; Chen et al. 2024, 2023a; Song et al. 2023b, 2024, a; Sun, Li, and Wang 2023; Sun et al. 2024a), there are pre-trained methodologies in place that have the capability to control the motion trajectories of objects within generated videos. VideoComposer (Wang et al. 2023c) employs a two-stage training strategy to incrementally incorporate temporal information and control signals. A motion condition encoder is proposed for training, designed to enhance the model’s ability to understand and integrate motion-specific information. DragNUMA (Yin et al. 2023) also employs a comparable training strategy and utilizes optical flow as a conditioning mechanism for trajectory modeling. MotionCtrl (Wang et al. 2023d) synthesizes the approaches of these two works and introduces a camera control module and a trajectory control module to further refine the management of movement within generated video content. Boximator (Wang et al. 2024a) proposes a novel self-tracking technique to simplify the learning of box-object relationships. MotionBooth (Wu et al. 2024) allows for controlling the movement of a specific object by fine-tuning a photo of that object. (Yu et al. 2024) enables the movement of a specified object in a user’s input photo to generate a video. Motion-I2V (Shi et al. 2024) leverages a two-stage training process, allowing users to alter the motion in the generated video by dragging. These models typically rely on extensive training on large-scale datasets such as WebVid-10M (Bain et al. 2021) and HD-VILA-100M (Xue et al. 2022) which lead to expensive training costs. Also, these models are typically constrained to operate on the specific models they were trained on, lacking the flexibility to interchange base video diffusion models. This limitation hinders their adaptability and limits the scope of their application to only those scenarios for which they were explicitly designed. Our proposed model stands out in that it can be applied to any base diffusion model without the necessity for further training. Recently, Trailblazer (Ma, Lewis, and Kleijn 2023) enhances attention on the box area in cross-attention with a zero-shot setting, similar to our work. In contrast, our operation on cross-attention is supervised by losses, providing better interpretability and effectiveness. Peekaboo (Jain et al. 2024) uses a masked attention module to achieve control without the need for training. FreeTraj (Qiu et al. 2024) utilizes frequency fusion to generate results whose trajectory is aligned with the given box. Although these methods exploit zero-shot settings in controllable video generalization, our proposed method emphasises on the importance of the initial noise and the control of spatial and temporal consistency, and achieves superior performance.
Preliminaries
@@
Video Diffusion Model: Video diffusion models are designed to produce high-quality and diverse videos, guided by text prompts. To save the computational costs, (Rombach et al. 2022) utilize a U-Net as a denoising model within a latent space, significantly reducing the computational load in terms of both time and space.
@@
Specifically, users can input a text condition $\mathbf{c}$ along with a sequence of rectangular boxes =$\mathcal{B}=\{(x_{1}^{f},y_{1}^{f}),(x_{2}^{f},y_{2}^{f})\}^{N_{f}}$, where $(x_{1}^{f},y_{1}^{f}),(x_{2}^{f},y_{2}^{f})$ are the upper left and lower right points of the box in frame $f$, $N_{f}$ is the total frame number. The boxes $\mathcal{B}$ correspond to the position of the moving object.
Figure 2: Overview of our Motion-Zero. The total pipeline is shown in (a). Given the box condition $\mathcal{B}$ and the prompt condition, we generate the prior latents $\mathbf{z}_{T}$ by our Initial Noise Prior Module (INPM) as shown on (b). At timestep $t$, $\mathbf{z}_{t}$ is firstly optimized to $\mathbf{z}_{t}^{\prime}$ by the Spatial Constraints (SC). Subsequently, $\mathbf{z}_{t}^{\prime}$ is passed to the UNet with Shift Temporal Attention Module (STAM) as demonstrated on (c). All the parameters of the video diffusion are frozen. $T_{1}$ represents the number of timesteps during which SC and STAM are applied, and $T_{2}$ denotes the number of timesteps where the original video diffusion process is utilized.
Methodology
Overview
@@
Existing trajectory control methods require large-scale video training data to optimize Eq.3, leading to a high computational cost. Differently, our work falls into the setting of zero-shot motion trajectory control. To this end, our proposed Motion-Zero framework operates entirely during the inference stage, thus eliminating the need for training and making it compatible with various pre-trained video diffusion models. The pipeline is shown in Fig.2 (a). In the following section, we provide a detailed presentation of our proposed Motion-Zero framework and its components. Firstly, we introduce a noise prior generation module to produce the initial noise for diffusion. Then, we describe the designed constraints and loss of the moving object position and spatial consistency. Finally, our shift temporal attention mechanism is presented to further improve the temporal consistency.
Initial Noise Prior Module
@@
According to the theory of DDIM Inversion (Ho, Jain, and Abbeel 2020), the initial noise has a significant impact on the final generated outcome. We introduce Initial Noise Prior Module (INPM) to leverage this property to provide a strong prior for the position of the moving object. Several steps are involved to integrate a moving object into a sequence of frames with a coherent prior as shown in Fig.2 (b). Firstly, given a prompt $\mathbf{c}$ and the boxes in the same location with $\{\mathcal{B}^{0}\}^{N_{f}}$, a meta video $V_{meta}$ is sampled, $\mathbf{z}^{*}\sim\mathcal{N}(0,\mathbf{I})$ as latent input, using the baseline video diffusion model and our proposed spatial constraints (introduced in the next section). This generated video has the target object staying at the location with $\{\mathcal{B}^{0}\}^{N_{f}}$ due to the spatial constraints. It is noteworthy that controlling the model to generate standing-still objects is much easier than controlling the movement of objects. Then, a video latent $\mathbf{z}_{meta}$ is generated based on $V_{meta}$ from Encoder $\mathcal{E}$. Once $\mathbf{z}_{meta}$ is prepared, we perform a DDIM Inversion to obtain the corresponding noise latent representation $\mathbf{z}_{I}$. We crop the latent representation within the box $\mathcal{B}^{0}$ for each frame, creating a sequence of latent patches containing the visual target. Subsequently, we use a local mixup operation (Zhang et al. 2017) to mix the latent patches and the initial noise $\mathbf{z}^{*}$ in the range of $\mathcal{B}^{f}$ frame by frame. Our INPM allows us to set a coherent prior in the corresponding object’s position in the initial noises. It also ensures that the animated object maintains consistency in appearance and movement across the video frames, without incurring additional computational costs during the generation process. Details are shown in Alg.1.
@@
Function: $\mathbf{Pix}(\mathbf{a},\mathbf{b})$ means the elements get from tensor $\mathbf{a}$ in the range of box $\mathbf{b}$
Output: Initial noise $\mathbf{z}_{T}$
@@
Spatial Constraints with Attention Map
The INPM alone is insufficient for precise manipulation of an object’s trajectory. To further improve the capability of control, we introduce Spatial Constraints (SC) deployed at each denoising step $t$ to optimize the intermediate latent representation $\mathbf{z}_{t}$. This optimization is crucial for enhancing the accuracy of the moving object position and preserving spatial consistency. Within the conditional denoising architecture (Rombach et al. 2022), cross-attention serves as the pivotal bridge that connects the text prompt with the content generated. During the denoising steps, conditioned on the prompt $\mathbf{c}$ and the intermediate features $\mathbf{z}_{t}$, the corresponding cross-attention map can be obtained as $\mathbf{A}$:
@@
where $\mathbf{Q},\mathbf{K}$ are the query and key with the $\mathbf{z}_{t}$ and $\mathbf{c}$, respectively. $\mathbf{W}_{Q},\mathbf{W}_{K}$ are two learnable matrices, which are frozen in our settings. Assuming a maximum number $N_{p}$ of prompt tokens $\{\mathbf{p}_{1},...,\mathbf{p}_{N_{p}}\}$ in condition $\mathbf{c}$, at time step $t$, it is feasible to derive $N_{p}$ cross-attention maps $\{\mathbf{A}_{1},...,\mathbf{A}_{N_{p}}\}$.
@@
$(W_{\mathbf{A}_{k}^{f}},H_{\mathbf{A}_{k}^{f}})=\frac{1}{\sum_{w,h}\mathbf{A}_{k,w,h}^{f}}(\sum_{w,h}w\cdot\mathbf{A}_{k,w,h}^{f},\sum_{w,h}h\cdot\mathbf{A}_{k,w,h}^{f})$,
${\mathcal{L}_{c}}^{f}=||(\frac{x_{1}^{f}+x_{2}^{f}}{2},\frac{y_{1}^{f}+y_{2}^{f}}{2})-(W_{\mathbf{A}_{k}^{f}},H_{\mathbf{A}_{k}^{f}})||_{1},$ (9)
where $\mathbf{A}_{k,w,h}^{f}$ is the value of $\mathbf{A}_{k}^{f}$ at the position of $(w,h)$, $(W_{\mathbf{A}_{k}^{f}},H_{\mathbf{A}_{k}^{f}})$ is the centroid position of $\mathbf{A}_{k}^{f}$ with the $k$-th token concept. After establishing robust control over the position of the object, we recognize unexpected variations in the object’s appearance due to the substantial extent of movement. To ensure appearance consistency across frames, we strive to maintain the uniformity of the $\mathbf{A}_{k}^{f}$ within the box. To this end, we introduce a similarity loss $\mathcal{L}_{s}$ as
[…]
where $\mathbf{Sim}(\cdot,\cdot)$ means the similarity of two elements and $\mathbf{Pix}()$ is a function getting corresponding elements of the map $\mathbf{A}^{f}$ within a box range $\mathcal{B}^{f}$. Cosine similarity is adopted here. At each timestep, the overall spatial constraints $\mathcal{L}_{sp}$ are formulated as follows:
${\mathcal{L}_{sp}}=\sum_{f}(\lambda_{i}{{\mathcal{L}_{i}}^{f}+\lambda_{o}\mathcal{L}_{o}}^{f}+\lambda_{c}{\mathcal{L}_{c}}^{f})+\lambda_{s}{\mathcal{L}_{s}},$ (11)
where $\lambda_{i},\lambda_{o},\lambda_{c},\lambda_{s}$ are hyper-parameters. By minimizing and calculating the gradient of the Eq.11, we can optimize our latent $\mathbf{z}_{t}$ in Eq.2 as follows:
@@
Shift Temporal Attention Mechanism
After applying the spatial constraints on attention maps, pre-trained video diffusion models still encounter difficulties in generating sequences of continuous actions. Within the temporal module of the diffusion process, the latent representation is reshaped into the following configuration:
$z_{(b\cdot H\cdot W+h\cdot W+w,f,c)}^{\prime}=z_{(b,f,c,h,w)}$ (13)
where $z^{\prime}$ represents the result of applying the rearrange operation to $z$. Within the temporal transformer, attention is focused on the same pixel across different frames. This leads to a scenario where, if the extent of motion is too large, the same position in different frames could undergo significant semantic changes, resulting in a lack of coherence in the generated dynamics.
To overcome this inconsistency, we propose a Shift Temporal Attention Mechanism (STAM) to improve the dynamics of the moving object in different frames. Specifically, we shift the elements of $\mathbf{z}^{f}$ inside the $\mathcal{B}^{f}$ range with the elements inside the $\mathcal{B}^{0}$ range and use the overlapped parts to fill in the vacated spaces, as shown in Fig.2 (c). Therefore, the subsequent frames within the box range can be aligned with the box range of the first frame. The steps are shown:
@@
where $\mathbf{Shift}(\cdot,a,b)$ is the shift operation, ${\mathbf{z}^{f}_{w}}$ means the shifted latent ${\mathbf{z}^{f}}$ of frame $f$ and ${\mathbf{z}_{w}}=[{\mathbf{z}^{1}_{w}},..,{\mathbf{z}^{f}_{w}}]$.By applying STAM to a $\mathbf{TemporalAttention}$ in the baseline video diffusion models, we can achieve coherence in the motion of moving objects without additional training, and without incurring extra computational costs during inference. Note that we do not use STAM in INPM as the motion of objects in the meta video $V_{meta}$ generated within INPM occurs at the same location. Therefore, there is no need to shift temporal attention to align the positions of moving objects.
Experiments
We evaluate the effectiveness of our method from both qualitative and quantitative perspectives. In our experiments, the default baseline is ZeroScope (Sterling 2023). ModelScope (Wang et al. 2023b) is also employed to show that our method can be applied to various video diffusion baselines. TrailBlazer (Ma, Lewis, and Kleijn 2023) and Peekaboo (Jain et al. 2024) are involved in experiments to demonstrate the control ability of our proposed method. Following TrailBlazer, 33 prompts containing different moving objects and motion patterns are employed as the evaluation dataset. For simple motions, we employed eight movement trajectories. For experiments involving complex trajectories, we utilized 17 randomly generated motion curves. Please refer to the appendix for specific prompts and trajectory parameters.
Implementation Details